OpenStudy (anonymous):

Help medal will be given!!!! Using the following equation, find the center and radius: x2 - 4x + y2 + 8y = -4 The center is located at (-2, -4), and the radius is 4. The center is located at (2, -4), and the radius is 4. The center is located at (-2, -4), and the radius is 16. The center is located at (2, -4), and the radius is 16.

OpenStudy (anonymous):

@pooja195 could you help plz

OpenStudy (welshfella):

convert to standard form by completing the square

OpenStudy (welshfella):

x2 - 4x + y2 + 8y = -4 = (x - 2)^2 - 4 + (y + 4)^2 - 16 = -4 = (x - 2)^2 + ( y + 4)^2 = 16

OpenStudy (anonymous):

so would it be B

OpenStudy (welshfella):

compare this with standard form (x - a)^2 + (y - b)^2 = r^2 where center = (a,b) and r = radius

OpenStudy (welshfella):

yes B

OpenStudy (anonymous):

thank you could you help me with another im stuck on

OpenStudy (welshfella):

ok

OpenStudy (anonymous):

ABCD is a quadrilateral inscribed in a circle:What equation can be used to solve for angle D?

OpenStudy (anonymous):

(2x + 16) + (6x - 4) = 90 (x + 16) - (x) = 90 (x + 16) + (2x + 16) = 180 (2x + 16) + (x) = 180

OpenStudy (welshfella):

opposite angles of a cyclic quadrilateral add up to 180 degrees

OpenStudy (welshfella):

what angle is opposite D?

OpenStudy (anonymous):

Is it B?

OpenStudy (welshfella):

no

OpenStudy (anonymous):

huh

OpenStudy (welshfella):

D + B = 180 - plug in the values of D and B

OpenStudy (anonymous):

(2x 16) + x =180

OpenStudy (welshfella):

yep

OpenStudy (anonymous):

thanks